# Mitigating the Accuracy-Robustness Trade-off via Multi-Teacher Adversarial Distillation

### v1

@@ 4.2 Ablation Study
Figure 6: The teachers’ information entropy of $T_{adv}$ and $T_{nat}$ with ResNet-18 student on CIFAR-10. The left is Baseline+NLB, and the right is MTARD (Baseline+NLB+EBB). NLB and EBB are abbreviations of Normalization Loss Balance and Entropy-based Balance. The x-axis means the training epochs (the first 40 epochs), and the y-axis is the entropy $T_{adv}$ and $T_{nat}$ in the training epoch x.
Figure 7: The result with different initial temperature settings in MTARD. The left is the results of different ranges for teachers’ temperatures. The right is the results of different initial values for teachers’ temperatures. The initial values of $\tau_{nat}$ and $\tau_{adv}$ for $T_{adv}$ and $T_{nat}$ are equal. All the results are the best checkpoint based on W-Robust Acc.
TABLE IV: Study with different $\beta$, the students are ResNet-18 trained on CIFAR-10 guided by ResNet-56 and WideResNet-34-10. All the results are the best checkpoint based on W-Robust Acc.
Attack
$\beta$ value
Clean
Robust
W-Robust
[…]
$\beta=0.5$
88.51%
50.24%
69.38%
$\beta=1$
88.20%
51.68%
69.94%
$\beta=4$
89.10%
50.01%
69.56%
$\beta=10$
88.63%
49.09%
68.86%
[…]
$\beta=0.5$
88.51%
53.43%
70.97%
$\beta=1$
88.20%
54.40%
71.30%
$\beta=4$
89.10%
53.04%
71.07%
$\beta=10$
88.63%
52.63%
70.63%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
$\beta=0.5$
88.51%
48.59%
68.55%
$\beta=1$
88.20%
49.88%
69.04%
$\beta=4$
89.10%
48.61%
68.86%
$\beta=10$
88.63%
48.16%
68.40%
@@
To better understand the impact of each component in our MTARD, we conduct a set of ablation studies. The baseline denotes using a clean teacher and a robust teacher to guide the student from clean and robust aspects, respectively, where the weight $w_{nat}$ and $w_{adv}$ are constant at 0.5. Baseline+NLB term denotes adding the Normalization Loss Balance (NLB) algorithm to dynamically adjust the weight $w_{nat}$ and $w_{adv}$ based on the Baseline. Baseline+EBB term denotes adding the Entropy-Based Balance (EBB) algorithm to dynamically adjust the $\tau_{nat}$ and $\tau_{adv}$ based on the Baseline. MTARD is the final version of our method. The performance is shown in Fig. 3. The change of total loss $L_{total}$ is shown in Fig. 4, and the change of relative loss $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ is shown in Fig. 5. Meanwhile, the change for the teacher’s information entropy is shown in Fig. 6.
@@
In Fig. 3, MTARD’s improvement brought by each component is remarkable, which shows its effectiveness. Multiple teachers positively affect the student model to learn with both clean and robust accuracy. However, it is not enough to mitigate the trade-off between accuracy and robustness without the Entropy-based Balance algorithm and the Normalization Loss Balance algorithm.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
[…]
SAT[25]
84.20%
64.84%
74.52%
56.16%
38.10%
47.13%
50.08%
33.40%
41.74%
TRADES[45]
83.00%
63.61%
73.31%
57.75%
38.20%
47.98%
48.45%
31.01%
39.73%
ARD[9]
84.11%
63.50%
73.81%
60.11%
39.53%
49.82%
53.22%
34.74%
43.98%
RSLAD[49]
83.99%
63.96%
73.98%
58.25%
39.93%
49.09%
48.78%
32.85%
40.82%
SCORE[26]
84.43%
64.34%
74.39%
56.40%
39.09%
47.75%
10.05%
8.74%
9.40%
MTARD(ours)
88.20%
65.29%
76.75%
65.08%
42.11%
53.60%
56.81%
36.65%
46.73%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
SAT [25]
84.20%
63.84%
74.02%
56.16%
39.42%
47.79%
50.08%
33.20%
41.63%
TRADES[45]
83.00%
62.83%
72.92%
57.75%
38.63%
48.19%
48.45%
30.72%
39.59%
ARD[9]
84.11%
62.86%
73.49%
60.11%
38.85%
49.48%
53.22%
33.32%
43.27%
RSLAD[49]
83.99%
63.05%
73.52%
58.25%
39.67%
48.96%
48.78%
32.09%
40.44%
SCORE[26]
84.43%
63.79%
74.11%
56.40%
40.02%
48.21%
10.05%
8.82%
9.44%
MTARD(ours)
88.20%
64.64%
76.42%
65.08%
41.35%
53.22%
56.81%
33.40%
45.11%
Square Attack[1]
SAT [25]
84.20%
72.48%
78.34%
56.16%
40.05%
48.11%
50.08%
38.72%
44.40%
TRADES[45]
83.00%
72.49%
77.75%
57.75%
42.51%
50.13%
48.45%
36.58%
42.52%
ARD[9]
84.11%
74.60%
79.36%
60.11%
47.20%
53.66%
53.22%
42.58%
47.90%
RSLAD[49]
83.99%
72.47%
78.23%
58.25%
45.32%
51.79%
48.78%
37.64%
43.21%
SCORE[26]
84.43%
74.68%
79.56%
56.40%
43.98%
50.34%
10.05%
8.67%
9.36%
MTARD(ours)
88.20%
79.82%
84.01%
65.08%
49.40%
57.24%
56.81%
44.46%
50.64%
TABLE VI: Black-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The result of three datasets is trained based on the MobileNet-v2. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
[…]
SAT [25]
83.87%
64.66%
74.27%
59.19%
40.70%
49.95%
49.03%
33.47%
41.25%
TRADES[45]
77.95%
61.04%
69.50%
55.41%
37.76%
46.59%
43.81%
28.35%
36.08%
ARD[9]
83.43%
63.28%
73.36%
60.45%
39.15%
49.80%
45.53%
30.73%
38.13%
RSLAD[49]
83.20%
64.33%
73.77%
59.01%
40.32%
49.67%
45.69%
31.20%
38.45%
SCORE[26]
82.32%
63.84%
73.08%
49.38%
40.00%
44.69%
28.27%
21.82%
25.05%
MTARD(ours)
88.70%
66.20%
77.45%
66.13%
42.67%
54.40%
52.98%
34.25%
43.62%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
SAT [25]
83.87%
64.24%
74.06%
59.19%
40.97%
50.08%
49.03%
33.13%
41.08%
TRADES[45]
77.95%
60.66%
69.31%
55.41%
38.02%
46.72%
43.81%
28.64%
36.23%
ARD[9]
83.43%
62.83%
73.13%
60.45%
38.53%
49.49%
45.53%
30.23%
37.88%
RSLAD[49]
83.20%
63.45%
73.33%
59.01%
39.92%
49.47%
45.69%
31.10%
38.40%
SCORE[26]
82.32%
63.03%
72.68%
49.38%
36.91%
43.15%
28.27%
22.19%
25.23%
MTARD(ours)
88.70%
65.73%
77.22%
66.13%
42.04%
54.09%
52.98%
33.50%
43.24%
Square Attack[1]
SAT [25]
83.87%
73.01%
78.44%
59.19%
44.51%
51.85%
49.03%
37.95%
43.49%
TRADES[45]
77.95%
67.43%
72.69%
55.41%
40.71%
48.06%
43.81%
32.39%
38.10%
ARD[9]
83.43%
73.26%
78.35%
60.45%
46.95%
53.70%
45.53%
34.60%
40.07%
RSLAD[49]
83.20%
73.07%
78.14%
59.01%
45.66%
52.34%
45.69%
35.18%
40.44%
SCORE[26]
82.32%
72.41%
77.37%
49.38%
37.94%
43.66%
28.27%
22.16%
25.22%
MTARD(ours)
88.70%
79.30%
84.00%
66.13%
50.83%
58.48%
52.98%
40.62%
46.80%
In Fig. 4, compared with RSLAD and Baseline, the Baseline+NLB’s training loss is less oscillating and can converge better. In Fig. 5, the gap between $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ can represent the difference of the knowledge proportion that students have not learned from the clean teacher and the robust teacher. In the final training period, the Baseline+NLB’s trade-off between $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ are tinier. The results demonstrate Normalization Loss Balance can indeed make the student have relatively equal knowledge learning speeds from different teachers.
In Fig. 6, we can see that in Baseline+NLB, the information entropy gap between the clean teacher and the robust teacher is very obvious, but after the adjustment of our Entropy-based Balance algorithm, the information entropy gap quickly narrows and remains relatively consistent, which proves the effectiveness of our Entropy-based Balance algorithm. Combined with the experimental results in Fig. 3, reducing the difference in teacher’s information entropy can indeed improve the performance of the student.
@@ 4.3 The discussion about hyper-parameters
Hyper-parameter $\tau$. The temperatures of different teachers are automatically adjusted by our Entropy-based algorithm until the information entropy is equal. However, the initial temperature setting still directly influences the amount of information entropy for both teachers, which has an inevitable impact on the final performance. Intuitively speaking, the temperature is supposed to change in a proper range to maintain the effectiveness of teachers’ knowledge (If the temperature is close to infinity, the predicted distribution is close to uniform distribution; If the temperature is close to 0, the predicted distribution is close to one-hot distribution). In order to explore the proper range for temperature, we select three ranges: [0.1, 1], [1, 10], and [10, 20] for teachers’ temperatures to change (The initial temperatures for different teachers are median of the ranges: 0.5, 5, 15, respectively). Based on the result shown in Fig. 7 (left), we select the [1, 10] as our initial setting. Meanwhile, to select the proper initial temperature value, we select initial $\tau_{nat}$ and $\tau_{adv}$ as 1, 2, 5, 8 in the range [1, 10]. Based on the result shown in Fig. 7 (right), we select 1 as our initial temperature value.
Hyper-parameter $\beta$. Here, we explore the role of $\beta$ in Normalization Loss Balance (Eq. (23) and Eq. (24)). We choose ResNet-56 and WideResNet-34-10 as the clean and robust teachers, respectively, and select ResNet-18 as the student on CIFAR-10. We test the results of different values $\beta$ for the training of the student. We select $\beta$ as 0.5, 1, 4, and 10 in our experiment. The result is presented in Table IV.
@@ 4.4 Adversarial Robustness Evaluation
White-box Robustness. The performances of ResNet-18 and MobileNet-v2 trained by our MTARD and other baseline methods under the white box attacks are shown in Table II and III for CIFAR-10, CIFAR-100, and Tiny-ImageNet. We select the best checkpoint of the baseline model and MTARD based on Weighted Robust Accuracy.
[…]
@@
Black-box Robustness. In addition, we also test MTARD and other methods against black-box attacks for ResNet-18 and MobileNet-v2 on CIFAR-10, CIFAR-100, and Tiny-ImageNet separately. We choose the transfer-based attack and query-based attack in our evaluation. We select the best checkpoint of all the trained models based on Weighted Robust Accuracy. The result of ResNet-18 is shown in Table V, while the result of MobileNet-v2 is shown in Table VI.
[…]
TABLE VII: The gradient estimation (GE) Attack [33] toward the ResNet-18 trained by MTARD on different datasets. The results are the robust Acc.
Attack
CIFAR-10
CIFAR-100
Tiny-ImageNet
[…]
54.40%
29.94%
24.94%
GE attack [33]
60.34%
32.53%
30.83%
@@
Excluding Obfuscated Gradients. Although some defense methods claim to resist adversarial attacks, their vulnerabilities are often exposed because they belong to “Obfuscated Gradients” [33, 2]. Athalye et al. [2] have confirmed that adversarial training [25] can effectively defend against the adaptive attack and do not belong to “Obfuscated Gradients”. Here we argue that MTARD is excluding the “Obfuscated Gradients” from the several aspects following [2] and [35]: (1) Our MTARD can effectively defend against strong attack: AutoAttack [6], which includes strong white-box attack: A-PGD and strong black-box attack: Square Attack [1]. (2) Strong white-box test attack (e.g. ${}_{\scriptscriptstyle\rm{\infty}}$[3]) have higher attack success rates than weak white-box test attack (e.g. FGSM [10]) in Table II and Table III. (3) White-box test attacks have higher attack success rates than Black-box test attacks (Comparing with Table II and Table V, Table III and Table VI). (4) To avoid insufficient attack rounds and to fall into the local optimal solution in the general attack methods, a gradient estimation attack method from [33] is directly applied to test MTARD. The result in Table VII shows MTARD can effectively resist the strong gradient estimation attack. Meanwhile, the black-box attack success rate is lower than white-box attack success rate, which further confirms MTARD is not “Obfuscated Gradients”.

### v3

@@ 4.2 Ablation Study
Figure 6: The teachers’ information entropy of $T_{adv}$ and $T_{nat}$ with ResNet-18 student on CIFAR-10. The left is Baseline+NLB, and the right is B-MTARD (Baseline+NLB+EBB). NLB and EBB are abbreviations of Normalization Loss Balance and Entropy-based Balance. The x-axis means the training epochs (the first 40 epochs), and the y-axis is the entropy $T_{adv}$ and $T_{nat}$ in the training epoch x.
@@
To better understand the impact of each component in our B-MTARD, we conduct a set of ablation studies. The baseline denotes using a clean teacher and a robust teacher to guide the student from both aspects, respectively, where the weight $w_{nat}$ and $w_{adv}$ are constant at 0.5. Baseline+NLB term denotes adding the Normalization Loss Balance (NLB) algorithm to dynamically adjust the weight $w_{nat}$ and $w_{adv}$ based on the Baseline, which is also the method in our ECCV version (MTARD) [59]. Baseline+EBB term denotes adding the Entropy-Based Balance (EBB) algorithm to dynamically adjust the $\tau_{nat}$ and $\tau_{adv}$ based on the Baseline, where the weight $w_{nat}$ and $w_{adv}$ are constant at 0.5. B-MTARD is the final version of our method, which applies both Normalization Loss Balance and Entropy-Based Balance algorithm. The performance is shown in Fig. 3. The change of total loss $L_{total}$ in the training process is shown in Fig. 4, and the change of relative loss $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ in the training process is shown in Fig. 5. The change for the teacher’s information entropy in the training process is shown in Fig. 6.
@@
In Fig. 3, B-MTARD’s improvement brought by each component is remarkable, which shows its effectiveness. Multiple teachers positively affect the student model to learn with both clean and robust accuracy. However, it is not enough to mitigate the trade-off between accuracy and robustness without the Entropy-based Balance algorithm and the Normalization Loss Balance algorithm, both of them can further enhance the W-Robust performance.
Baseline+NLB outperforms Baseline by 0.87%, 0.56%, and 0.6% against the attack of PGDsat, PGDtrades, and CW in the metric of W-robust Acc, while B-MTARD outperforms Baseline+EBB by 0.33%, 0.23%, and 0.15% against the attack of PGDsat, PGDtrades, and CW in the metric of W-robust Acc. The improvement shows the effectiveness of our Normalization Loss Balance Algorithm.
In addition, B-MTARD has an obvious improvement compared with our ECCV version (Baseline + NLB): Our B-MTARD improves W-Robust Acc by 0.89%, 0.82%, and 0.94% compared with Baseline + NLB against the attack of PGDsat, PGDtrades, and CW. The result denotes the necessity and effectiveness of our incremental work (the Entropy-based Balance algorithm) in this version.
In Fig. 4, compared with RSLAD and Baseline, the Baseline+NLB’s training loss is less oscillating and can converge better. In Fig. 5, the gap between $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ can represent the difference of the knowledge proportion that students have not learned from the clean teacher and the robust teacher. The Baseline+NLB’s trade-off between $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ are tinier. The results demonstrate Normalization Loss Balance can indeed make the student have relatively equal knowledge learning speeds from different teachers.
In Fig. 6, we can see that in Baseline+NLB, the information entropy gap between the clean teacher and the robust teacher is very obvious, but after the adjustment of our Entropy-based Balance algorithm, the information entropy gap quickly narrows and remains relatively consistent, and this phenomenon can denote the knowledge scale gap between clean teacher and robust teacher reduces, which proves the effectiveness of our Entropy-based Balance algorithm. Combined with the experimental results in Fig. 3, reducing the difference in teacher’s information entropy can indeed improve the performance of the student.
@@ 4.3 The discussion about hyper-parameters
Hyper-parameter $\tau$. The temperatures of different teachers are automatically adjusted by our Entropy-based algorithm until the information entropy is equal. However, the initial temperature setting still directly influences the amount of information entropy for both teachers and determines the prior adjustment scope, and a proper initial temperature can ensure temperature changes within a reasonable range to maintain the effectiveness of teachers’ knowledge. To select the proper initial temperature value, we select initial $\tau_{nat}$ and $\tau_{adv}$ as 1, 2, 5, 8, while all other experimental settings remain unchanged. Based on the results shown in Table VII, we select 1 as our initial temperature value.
TABLE VII: Study with different initial $\tau$. $\beta$ is fixed to 1, the students are ResNet-18 trained on CIFAR-10 guided by ResNet-56 and WideResNet-34-10.
Attack initial $\tau$ Clean Robust W-Robust
PGDsat[31] $\tau=1$ 88.20% 51.68% 69.94%
$\tau=2$ 88.22% 51.43% 69.83%
$\tau=5$ 89.27% 47.89% 68.58%
$\tau=8$ 88.64% 46.97% 67.81%
PGDtrades[55] $\tau=1$ 88.20% 54.40% 71.30%
$\tau=2$ 88.22% 54.13% 71.18%
$\tau=5$ 89.27% 50.90% 70.09%
$\tau=8$ 88.64% 49.50% 69.07%
CW[4] $\tau=1$ 88.20% 49.88% 69.04%
$\tau=2$ 88.22% 49.79% 69.01%
$\tau=5$ 89.27% 43.99% 66.63%
$\tau=8$ 88.64% 42.45% 65.55%
TABLE VIII: Study with different $\beta$. Initial $\tau$ is fixed to 1, the students are ResNet-18 trained on CIFAR-10 guided by ResNet-56 and WideResNet-34-10.
Attack $\beta$ Clean Robust W-Robust
PGDsat[31] $\beta=0.5$ 88.51% 50.24% 69.38%
$\beta=1$ 88.20% 51.68% 69.94%
$\beta=4$ 89.10% 50.01% 69.56%
$\beta=10$ 88.63% 49.09% 68.86%
PGDtrades[55] $\beta=0.5$ 88.51% 53.43% 70.97%
$\beta=1$ 88.20% 54.40% 71.30%
$\beta=4$ 89.10% 53.04% 71.07%
$\beta=10$ 88.63% 52.63% 70.63%
CW[4] $\beta=0.5$ 88.51% 48.59% 68.55%
$\beta=1$ 88.20% 49.88% 69.04%
$\beta=4$ 89.10% 48.61% 68.86%
$\beta=10$ 88.63% 48.16% 68.40%
Hyper-parameter $\beta$. Here, we explore the role of $\beta$ in Normalization Loss Balance (Eq. (17) and Eq. (18)). We test the results of different values $\beta$ for the training of the student. We select $\beta$ as 0.5, 1, 4, and 10 in our experiment, while all other experimental settings remain unchanged. The results are presented in Table VIII.
@@ 4.4 Adversarial Robustness Evaluation
White-box Robustness. The performances of ResNet-18 and MobileNet-v2 trained by our B-MTARD and other baseline methods under the white box attacks are shown in Table III and IV for CIFAR-10, CIFAR-100, and Tiny-ImageNet.
The results in Table III and IV demonstrate that B-MTARD achieves state-of-the-art W-Robust Acc on CIFAR-10, CIFAR-100, and Tiny-ImageNet. For ResNet-18, B-MTARD improves W-Robust Acc by 1.44%, 2.20%, and 1.41% compared with the best baseline method against the PGDtrades on CIFAR-10, CIFAR-100, and Tiny-ImageNet; B-MTARD improves W-Robust Acc by 0.79%, 2.76%, and 1.54% compared with the best baseline method against the PGDtrades on CIFAR-10, CIFAR-100, and Tiny-ImageNet. Moreover, B-MTARD shows relevant superiority against FGSM, PGDsat, CW, and AutoAttack compared with other methods. The results show that the overall performance of B-MTARD has different degrees of advantage whether compared with the other baseline methods. Meanwhile, B-MTARD also shows the corresponding superiority compared with our conference version (MTARD).
@@
Black-box Robustness. In addition, we also test B-MTARD and other methods against black-box attacks for ResNet-18 and MobileNet-v2 on CIFAR-10, CIFAR-100, and Tiny-ImageNet separately. We choose the transfer-based attack and query-based attack in our evaluation. We select the best checkpoint of all the trained models based on W-Robust Acc. The results of ResNet-18 are shown in Table V, while the results of MobileNet-v2 are shown in Table VI.
From the result, B-MTARD trained models achieve the best W-Robust Acc against all three black-box attacks compared to other models. Under the Square Attack, B-MTARD trained ResNet-18 improves W-Robust Acc by 4.45%, 3.58%, and 2.74% on CIFAR-10, CIFAR-100, and Tiny-ImageNet compared to the second-best method; Moreover, B-MTARD brings 5.56%, 4.78%, and 3.31% improvements to MobileNet-v2 on CIFAR-10, CIFAR-100, and Tiny-ImageNet. In addition, B-MTARD has different margins in defending against query-based attack of RayS and transfer attacks of PGDtrades and CW, which shows the superior performance of B-MTARD in black-box robustness.
@@
Excluding Obfuscated Gradients. Although some defense methods claim to resist adversarial attacks, their vulnerabilities are often exposed because they belong to “Obfuscated Gradients” [40, 3]. Athalye et al. [3] have confirmed that Adversarial Training [31] can effectively defend against the adaptive attack and do not belong to “Obfuscated Gradients”. Here we argue that B-MTARD is excluding the “Obfuscated Gradients” from the several aspects following [3] and [42]: (1) Our B-MTARD can effectively defend against strong attack: AutoAttack [10], which includes strong white-box attack: A-PGD and strong black-box attack: Square Attack [2]. (2) Strong white-box test attacks (e.g., CW[4]) have higher attack success rates than weak white-box test attacks (e.g., FGSM [14]) in Table III and Table IV. (3) White-box test attacks have higher attack success rates than Black-box test attacks (Comparing with Table III and Table V, Table IV and Table VI). (4) To avoid insufficient attack rounds and to fall into the local optimal solution in the general attack methods, a gradient estimation attack method from [40] is directly applied to test B-MTARD. The results in Table IX show B-MTARD can effectively resist the strong gradient estimation attack. Meanwhile, the black-box attack success rate is lower than the white-box attack, which further confirms B-MTARD is not “Obfuscated Gradients”.
TABLE IX: The gradient estimation (GE) Attack [40] toward the ResNet-18 trained by B-MTARD on different datasets. The results are the robust Acc.
Attack CIFAR-10 CIFAR-100 Tiny-ImageNet
Base PGDtrades[55] 54.40% 29.94% 24.94%
GE attack [40] 60.34% 32.53% 30.83%
4.5 Comparison with other Multi-teacher Distillation
In order to further prove our effectiveness, we apply the representative multi-teacher distillation method [51] into the framework of Adversarial Training to compare with our method. Here we apply the same adversarial teacher and clean teacher with our B-MTARD for [51]. We apply the weighted average of different teacher output logits as the guidance for both the adversarial examples and clean examples to train the student model (ResNet-18) on CIFAR-10. The results are shown in Table X.
TABLE X: The Comparison with other Multi-teacher Distillation. The results are trained on CIFAR-10 of ResNet-18. The results are on W-Robust Acc.
Attack FGSM PGDsat PGDtrades CW AA
LMTN[51] 72.56% 68.50% 69.62% 68.12% 67.09%
B-MTARD 74.81% 69.94% 71.30% 69.04% 67.82%
The results show the effectiveness of our proposed method. Our B-MTARD outperforms the LMTN by 2.25%, 1.44%, 1.68%, 0.92%, and 0.73% against the attack of FGSM, PGDsat, PGDtrades, CW, and AA in the metric of W-robust Acc. Actually, just as the discussion in related work, since the training of clean samples and adversarial samples will interfere with each other and lead to the existence of the trade-off phenomenon, how to choose an appropriate balance adjustment strategy to eliminate the existence of trade-off as much as possible is a difficult optimization challenge. Directly applying previous multi-teacher optimization methods has shortcomings in this new task scenario, which has been proven by experimental results. For that, we propose the Entropy-Based Balance algorithm and Normalization Loss Balance algorithm to balance the two optimization directions from the perspective of teacher and student, respectively, which do provide a balance in the conflicting multi-teacher distillation. The results further show the necessity and superiority of those two algorithms.